ESTIMATION OF ROTATIONAL TEMPERATURE OF ^{121}Sb^{16}O MOLECULE M. B. SURESHKUMAR, N. R. SHAH and \underline{ASHISH JAITUNI}, Department of Physics, Faculty of Science, The M. S. University of Baroda, Vadodara - 390002, India.